A Bibliography of Publications of Frank Stenger

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 August 2025 Version 2.01

Abstract

This bibliography records publications of Frank Stenger.

Title word cross-reference

3 + 1 [159]. 3 + 1^{1} [160]. R^{n} [39, 45]. Δu = u − u^{3} [40]. Γ [106]. H_{p} [64, 72, 74, 54, 55, 73]. I^{(−1)} [179]. LU [192]. n [43]. R^{3} × [0, T] [161]. |x|^{α} [98].

-space [43].

1-2 [201]. 11 [19]. 12 [53]. 14th [198]. 1973 [198]. 1980 [200]. 1983 [202]. 1988 [208]. 1990 [211]. 1993 [214].

2000 [218].

614 [72]. 65th [210].

80th [225]. 85th [221].

A. [29, 35]. acceleration [98]. Accuracy [187]. Acoustic [70]. Acoustical [203, 199, 207, 203]. Adaptive [66]. adsorption [152]. Algorithm [50, 66, 31, 43, 60, 61, 175, 72]. Algorithms [68, 69, 64]. AMS [219]. analogue [47]. Analysis [224, 84, 220, 137, 210, 225, 216, 204, 219]. Analytic [176, 111, 112, 42, 194, 75, 213]. and/or [79]. Angles [41, 46]. Anisotropic [143]. Any [192]. Appl [53]. Applicable [224]. Application [191, 98]. applications [215, 198, 212, 129]. applied [98]. applying [139]. Approach [182, 85, 158, 138]. Approximate [35, 42, 164, 40, 36, 27, 31, 32, 77, 93, 108, 138, 141, 33]. Approximating [163, 75]. Approximation [120, 224, 202, 114, 17, 95, 96, 104, 106, 107, 154, 194, 221, 98, 94, 115, 121, 145, 22, 81, 135, 140, 149, 214]. Approximations [78, 51, 157, 165, 167, 158, 56, 54, 55, 148]. April [206]. Arbitrary [1]. arc [86]. arcs [135]. Arising [183]. Asymptotic [1, 17, 65, 210, 2, 3, 6, 7, 18, 30]. Atlanta [201]. Audits [187]. Aug [198]. August [208, 211]. Austin [197]. Axially [50, 60]. Axially-Symmetric [50].

B [223]. band [92]. band- [92]. Based [176, 111, 112, 34, 53, 58, 61, 62, 65, 213, 140, 148]. basis [69]. Beach [206]. Beam [142]. BEM [137]. between [44]. bi [134]. bi-material [134]. biennial [198]. biharmonic [144]. Birthday [210, 221, 225]. Bisecting [41, 46]. Bisection [67]. bisections [47]. Body [50, 60]. Bolted [137]. Boltzmann [110]. Bonded [137]. Bonded/Bolted [137]. Book [120, 176, 121, 12, 21, 28, 33, 88, 95, 100, 119, 129]. Books [123]. Bound [41, 46, 190]. Boundary [183, 50, 91, 57, 60, 134, 129]. Bounds [49, 1, 5, 9, 2, 3, 6, 7, 8]. Bowers [100, 105]. Bozeman [208, 211]. Bremen [200]. Bubnov [38]. Burgers [108, 195]. Butzer [221].

C [21, 24]. Cairo [216]. Calculation [33, 35]. Calculations [78]. calculus [150]. California [205, 206]. Can [26]. Canada [207]. Canadian [198]. Canas [129]. Cardinal [56, 20, 51, 62]. Cardinal-type [56]. case [2, 3, 6, 7]. Cathedral [205]. Cauchy [86, 44, 135]. Cauchy-type [135]. Certain [19, 17]. Chakravarti [21, 24]. Characteristic [42]. Characterization [199]. Chemistry [159, 160, 222, 223]. Chi [215]. circumventing [117]. City [215]. classical [117]. CMFT [217]. Collocating [116]. Collocation [183, 195, 132, 139]. comb [153]. Combined [84]. Comments [172]. Completely [188]. Complex [90, 224, 86]. complexity [92, 88, 89]. Composite [137, 143]. Computation [181, 114, 214, 120, 208, 211, 121]. Computational [89, 159, 160, 210, 93, 217, 88]. computations [15]. Computerized [206]. Computing [39, 45, 50, 119, 128, 158, 118, 156]. Concepts [222, 223]. Condition [50, 115, 60]. Conduction [182]. Conference [197, 202, 216, 218, 217, 210, 208, 211, 214]. Conformal [130]. Congress [198]. conjecture [179]. conjectures [180]. Connection [44]. Constant [157]. Constructed [41, 46]. Construction [4, 10, 194]. Constructive [22, 85]. Contributions [204]. control [208, 211, 198, 93].

Converge [26]. Convergence [98, 38, 54, 55]. Convolution [143, 36, 32, 108, 117, 128]. convolution-type [32]. convolutions [116]. corner [171]. corners [126]. Corrected [66]. Coupled [49, 195]. Cracks [137, 143]. Critical [169, 170]. Cyprus [217].

Data [79, 84, 149]. Davis [12, 14]. December [202, 214]. decreasing [190]. dedicated [178]. definite [30]. Degree [19, 39, 43, 45, 173, 175]. derivative [149]. derivatives [56]. Descriptions [123]. Detection [59]. Determinants [188]. Developments [70, 212]. Differential [197, 49, 1, 191, 9, 100, 101, 105, 204, 2, 3, 6, 7, 18, 77]. Diffraction [66]. diffusion [152]. Dimensional [137, 86, 47, 23, 25, 134]. discontinuous [115]. Discrete [137, 143]. Distinct [52, 2, 6]. Domains [125, 126, 127]. double [126]. double-layer [126]. due [91, 60].

Editors [178]. Effect [142]. Egypt [216]. Eigenfunctions [184]. eigenvalue [2, 6]. Eigenvalues [181, 184]. elasticity [134]. Election [187]. Electromagnetic [50, 60, 97]. Electronic [218]. element [129]. Elementary [168]. Elements [142]. energy [92]. energy-limited [92]. Engineering [206]. Ensure [187]. Equation [183, 59, 184, 82, 125, 162, 86, 110, 40, 36, 126, 115, 127, 58, 61, 108, 118]. Equations [200, 197, 67, 68, 78, 84, 189, 49, 1, 191, 9, 101, 102, 124, 161, 195, 47, 63, 132, 139, 133, 2, 3, 6, 7, 18, 27, 31, 32, 77, 123, 204, 100, 105]. equiconvergence [87]. Erratum [23]. Error [49, 1, 2, 3, 6, 7, 8, 9, 174, 190, 5, 38]. estimate [174]. estimates [55]. evaluation [8]. Exact [82, 70, 78, 84]. expansions [30]. Explicit [59, 81, 93]. Extensions [13]. Extrapolation [66, 76]. extremal [37].

F. [172]. Factorization [192]. Falkner [183]. Fast [82, 68, 83, 69]. Festschrift [225, 214, 221]. Field [78, 50]. filter [94]. finite [23, 25]. first [139, 18]. Florida [202]. flow [91]. focusing [153]. Fokker [162]. form [134]. formula [174, 190, 34, 53, 173]. formulae [8, 34, 53, 28, 29]. Formulas [19, 26, 194, 4, 10, 172]. FORTRAN [72, 73]. forward [118]. Fourier [44, 147]. Fractional [150, 152, 162, 191, 158]. Frame [142]. Frames [142]. Francisco [205]. Frank [178, 120, 176, 121, 112, 225, 185, 177, 19, 111, 222, 223, 155, 210]. Frederico [129]. Fredholm [188, 189]. free [91]. Frequencies [82]. Frequency [84, 66, 153, 65, 76]. Fully [19, 4, 10]. Function [83, 20, 42, 169, 170, 56, 217, 51, 75, 76, 140, 147, 149]. Functional [133]. Functions [194, 94, 190, 133, 22, 62, 75, 213, 140, 163, 176, 111, 112, 95, 96]. fundamentals [129].

G [88, 89]. Galerkin [71, 38, 57]. Gander [119]. Gauss [5, 8]. Gauss-type [5, 8]. Gaussian [182, 145]. Gautschi [214]. general [3, 7]. generated [153]. geophysics [215]. Georgia [201]. Ghizetti [28, 29]. given [86, 36]. gradient [85]. Green [140]. Guest [178].

H [33, 35]. Hackbusch [123, 124]. Halifax [207]. Handbook [151]. Harmonic [125, 127, 140]. Harmonic-Sinc [125, 127]. harmonics [153]. Harris [222, 223]. having [1]. Heat [182, 115]. held [202, 203]. Helmholtz [59, 82, 11, 68, 84, 58, 61, 118]. Hilbert [48]. Hill [205]. History [155]. Honor [210, 221, 225, 214]. Hopf [27, 31]. Hotel [201, 205]. Hrebícek [119]. Hybrid [137]. Hypothesis [166, 168].

Identities [165, 167]. IEEE [201, 205]. II [211, 92, 7]. III [69]. Image [196, 219]. Imaging [207, 206, 199, 203, 219]. Impedance [50, 60]. Improved [154]. impulse [94]. incompressible [132]. Indefinite [165, 167, 99, 131, 164]. industry [215]. Infinite [125, 127, 190]. Influence [186]. initial [115, 99]. inner [22]. Insonification [59, 58]. Inst [53]. Integral [189, 102, 86, 126, 27, 31, 32, 134, 123, 124]. Integrals [17, 24, 190, 8, 23, 25, 30, 135, 23, 33, 35, 21]. Integration [19, 72, 73, 34, 53, 164, 165, 167, 212, 4, 10, 99, 131, 12, 14]. Integrity [187]. Interaction [219]. Internal [78]. International [215, 216, 207, 203, 206]. Interpolant [186]. Interpolation [202]. Intersociety [218]. Interview [177]. Inverse [215, 182, 79, 82, 68, 69, 70, 78, 84, 219, 63, 215, 219]. Inversion [59, 90, 58, 61, 65, 97, 118]. Irregular [1]. Issue [178]. iteration [133]. Iterative [68, 78]. ITherm [218]. IVP [131].

J [12, 14, 53, 100, 210]. J. [105]. January [215, 216, 219]. Jírí [119]. Joints [137, 143]. Jose [129]. July [207]. Jumps [186]. junction [153].

K. [89]. Kalaba [44]. Kelvin [11]. Kingdom [202]. Kowalski [120, 121]. Kronecker [13]. Krzysztof [120, 121].

L [100, 105]. Laminated [143]. Laplace [90, 125, 127]. Large [195, 18]. laser [153]. Layer [183, 126]. Lebesgue [157]. lectures [209]. Lévy [184]. limitations [117]. Limited [82, 92]. Line [169, 170]. Linear [13, 18, 81]. Linewidth [153]. Located [169, 170]. locked [153]. Longest [41, 46]. Louisiana [219]. Lower [41, 46, 55]. Lubbock [209]. Lund [100, 105].

MAPLE [119]. mapping [39, 43, 45]. maps [130]. Marek [120, 121]. Marriott [201]. material [134]. Math [53]. Mathematical [216, 198, 103, 194, 204, 222, 223, 216]. MATLAB [119]. Matrices [122]. Matrix [192, 179]. May [218]. medical [219, 128]. medicine [215]. melting [85]. Memory [224]. Method [183, 67, 195, 86, 47, 69, 126, 132, 139, 38, 57, 129, 131]. Methods [181, 184, 71, 176, 111, 112, 100, 105, 109, 151, 159, 160, 193, 90, 150, 225, 179, 217, 80, 30, 62, 75, 77, 93, 213, 113, 122, 130, 136, 140, 141, 146, 156]. microscope [110]. microwave [153]. Minh [215]. minimum [54].

Minneapolis [203]. MN [203]. mode [153]. mode-locked [153]. model [110]. Modeling [220]. modified [190]. Moment [142, 71, 69]. Monotonic [188]. Montana [208, 211]. Movable [16]. Multi [191]. Multi-Order [191]. Multidimensional [206]. Multigrid [80]. Multigrid-sinc [80]. Multiple [82, 33, 35, 69]. multiplier [145]. Multiscale [220].

Navier [132, 139, 161]. nearly [81]. Nevada [218]. Newport [206]. Newton [131]. Nicosia [217]. Noise [66]. Nonlinear [200, 110, 67, 49, 85, 40, 47, 139]. norm [54]. Novel [134]. November [201]. NS [207]. Number [195, 23, 25]. Numerical [183, 197, 19, 111, 72, 14, 62, 213, 113, 124, 151, 164, 225, 86, 48, 4, 10, 204, 123, 136, 200, 212, 176, 112, 12]. Nyström [86].

obtaining [30]. October [197, 203, 201, 205, 217]. ODE [131]. ODE-IVP-PACK [131]. Olver [210]. one [86, 23, 25, 65]. one-dimensional [86, 23]. Ontario [198]. open [97]. Operators [13, 165, 167, 158]. Optimal [92, 64, 74, 54, 81, 198]. Optimization [194]. Order [1, 191, 18]. Ordinary [197]. Orleans [219]. orthogonal [180]. Ossicini [28, 29]. Outcome [187]. Overcome [186]. Overview [181].

P [21, 95, 96]. P. [24, 96]. PACK [131]. Parabolic [49]. parameter [18]. Paris [129]. Part [69, 223]. partial [204, 77]. Paul [221]. PDE [138, 141, 146]. penetration [91]. Performance [82]. Periodic [148]. Perspectives [221]. Petrushev [95, 96]. phase [85]. Phenomena [218]. Philip [12, 14, 14]. Physics [206, 159, 160, 222, 223]. planar [126]. Planck [162]. plane [86, 58]. point [85]. Poisson [110, 83]. poles [81]. Poly [195]. Poly-Sinc [195]. Polyhedra [193]. Polynomial [75, 104, 149]. polynomials [180, 37]. Popov [95, 96]. preassigned [81]. preface [178]. Preliminaries [103]. prescribed [37]. presented [218]. Problem [182, 70, 85, 91, 48, 37]. Problems [125, 119, 164, 215, 11, 127, 219, 57, 97, 99, 128, 134, 144, 156]. Proceedings [200, 202, 207, 203, 215, 208, 211, 198, 217, 214, 197]. Processing [206, 216, 117]. Product [13]. Professor [178]. Progress [224]. Projection [189]. Proof [179, 166, 168]. proofs [22]. Publications [185]. Purdue [214].

Q. [224]. Quadrature [16, 26, 29, 100, 105, 174, 126, 190, 172, 64, 28]. Quadratures [74, 5].

Rabinowitz [12, 14]. Radial [140]. Rahman [224]. Rank [1]. rapidly [190]. rate [55]. Rational [94, 76, 96, 75, 81, 202, 95]. Rayleigh [11]. Real [95, 96]. recovery [92]. Reduced [142]. reduction [23, 25]. Reflection [79, 84]. regularization [90]. Related [109, 133, 156, 163]. remark [175]. Remarks [53]. repeated [8]. representation [44]. Research [155, 97]. response [36]. Retrospect [20]. Review [120, 19, 176, 111, 112, 12, 14, 21, 24, 28, 29, 33, 35, 88, 89, 95, 96, 100, 105, 124].

Reviews [121, 119, 123, 129]. Reynolds [195]. Riccati [68, 78, 84]. Riemann [163, 166, 168, 169, 170]. Rivlin [87]. RYTOV [78].

Sampling [141, 221]. San [205]. scanning [110]. Scattered [50]. Scattering [79, 82, 68, 70, 78, 84, 69, 60]. scholar [209]. Schrödinger [159, 184, 160]. Science [196]. Scientific [119]. Second [1, 211]. Second-Order [1]. Section [142]. Selected [120, 114, 196, 121]. Semi [125, 127]. Semi-Infinite [125, 127]. seminar [198]. Separation [146]. series [147]. Session [219]. Seventh [218]. Side [41, 46]. Signal [220, 216, 117]. signals [92]. Sikorski [120, 121]. Simple [169, 170]. simplification [173]. Sinc [181, 182, 183, 158, 184, 225, 186, 125, 111, 115, 189, 191, 112, 145, 105, 106, 107, 109, 143, 146, 151, 156, 157, 193, 195, 69, 126, 127, 80, 62, 75, 134, 148, 150, 86, 98, 179, 83, 139, 57, 77, 99, 213, 108, 117, 118, 122, 128, 130, 131, 135, 136, 137, 144, 147, 149, 100]. Sinc-approximations [158]. Sinc-Collocation [183]. Sinc-Gaussian [182]. Sinc-Methods [181]. Sinc-Related [109]. Sine [176, 132]. Single [84]. singular [86]. Singularities [125, 16, 127]. Singularity [1]. Skan [183]. Slowly [26]. smooth [86]. soft [63]. software [212]. Solution [200, 183, 197, 125, 143, 164, 86, 40, 48, 204, 115, 127, 132, 18, 27, 31, 32, 57, 77, 108, 118, 134, 138, 141, 144, 146]. Solutions [79, 82, 68, 78, 84, 189, 49, 1, 9, 63, 69, 2, 3, 6, 7, 128, 156]. Solver [83]. Solving [182, 70, 119, 195, 47]. Some [97, 155, 172, 117]. sonic [97]. Sound [59]. source [69]. space [43]. Spaces [73, 74, 64]. Special [219]. sphere [171]. Spherical [59, 58]. stability [11]. States [202]. Stefan [85]. Stenger [225, 185, 174, 178, 120, 177, 19, 180, 179, 176, 111, 190, 121, 172, 112, 155, 173, 175]. step [94, 139]. Stiffness [142]. Stokes [132, 139, 161]. Strip [107, 169, 170]. Strongly [49]. Stroud [33, 35]. Subroutine [72, 73]. subsets [52]. Summary [136]. sums [52, 21, 24]. Symmetric [19, 50, 4, 10, 60]. Symposium [207, 203, 201, 205]. system [44]. Systems [67, 218, 49].

Tables [123]. Tabulation [19]. Tampa [202]. Tech [209]. technology [215]. Texas [197, 209]. Their [184, 167, 165]. theorem [87]. theoretical [190]. Theory [183, 224, 123, 124, 221, 180, 198, 217, 93]. Thermal [218]. Thermomechanical [218]. Third [217]. thirteenth [203]. Three [137]. tissue [63]. Tolerant [66]. Tomography [71, 66, 58, 61, 76]. tool [117]. Topics [120, 114, 196, 121]. topological [39, 43, 45, 173, 175]. Transform [30, 90]. Transformations [154, 113]. transforms [44]. Transmission [79, 84, 58]. trap [15]. trapezoidal [34, 53]. Treatment [124, 123]. Triangles [41, 46]. tribute [222, 223]. tunneling [110, 153]. two [85, 47, 18, 23, 25, 134]. two-dimensional [47, 134]. two-phase [85]. TX [209]. type [11, 56, 5, 8, 32, 135].

Ultra [59]. Ultra-Sound [59]. ultrafast [153]. Ultrasonic [58, 61, 65, 76].

Ultrasonics [201, 205]. Ultrasound [79, 71, 66]. unified [138]. unilateral [190]. Uniqueness [49]. United [202]. University [197, 216, 198, 209]. Upper [55]. USA [218, 209]. Use [154]. Using [181, 183, 184, 119, 82, 132, 65, 140].

V [95, 96]. value [91, 57, 99]. variable [90]. variables [146]. Vegas [218]. via [51, 113, 128, 130, 131, 146, 147, 154, 164]. Views [82]. viscoelastic [156]. Viscoplastic [91]. Visiting [209]. Visualization [199].

W [210]. Wagner [88, 89]. Walsh [87]. Walter [214, 119]. Wave [63, 68, 78, 84, 58]. wavelets [216]. Ways [186]. Wechsung [88, 89]. Western [198]. which [42]. Whittaker [20, 51, 62]. Wiener [27, 31]. without [118]. Wolfgang [123, 124]. Workshop [215, 206].

zero [37]. Zeros [169, 170]. Zeta [169, 170, 147, 163].

References

Olver:1965:EBAb

[1] F. W. J. Olver and F. Stenger. Error bounds for asymptotic solutions of second-order differential equations having an irregular singularity of arbitrary rank. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):244–249, ???? 1965. CODEN ???? ISSN 0887-459X (print), 2168-3581 (electronic).

URL http://links.jstor.org/sici?sici=0887-459X(1965)2:2<244:
EBFASO>2.0.CO%3B2-N.

Stenger:1965:EBAa

[2] Frank Stenger. Error bounds for asymptotic solutions of differential equations. 1, The distinct eigenvalue case. Technical Report 2, Department of Computing Science, University of Alberta, Edmonton, AB, Canada, 1965.

34 pp. URL http://web.sirsitest.library.ualberta.ca/uhtbin/
cgisirsi/QWzmKzdaVV/UAARCHIVES/295120075/9?first_hit=1&last_
hit=20&form_type=&VIEW%5E3.x=49&VIEW%5E1.y=10.

Stenger:1965:EBAb

[3] Frank Stenger. Error bounds for asymptotic solutions of differential equations. 2, The general case. Technical Report 3, Department of Computing Science, University of Alberta, Edmonton, AB, Canada, 1965.

44 pp. URL http://web.sirsitest.library.ualberta.ca/uhtbin/
cgisirsi/3DeRFr4iWs/UAARCHIVES/295120075/9?first_hit=1&last_
hit=20&form_type=&VIEW%5E4.x=49&VIEW%5E1.y=10.

McNamee:1966:CFS

[4] J. McNamee and F. Stenger. Construction of fully symmetric numerical integration formulas. Technical Report 4, Department of Computing Science, University of Alberta, Edmonton, AB, Canada, 1966. 32 pp.

Stenger:1966:BEG

[5] F. Stenger. Bounds on the error of Gauss-type quadratures. Numerische Mathematik, 8(2):150–160, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1966:EBAa

[6] Frank Stenger. Error bounds for asymptotic solutions of differential equations. I: The distinct eigenvalue case. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 70(3):167–186, July/September 1966. CODEN JNBBAU. ISSN 0022-4340 (print), 2376-5283 (electronic).

Stenger:1966:EBAb

[7] Frank Stenger. Error bounds for asymptotic solutions of differential equations. II: The general case. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 70(3):187– 210, May 1966. CODEN JNBBAU. ISSN 0022-4340 (print), 2376-5283 (electronic).

Stenger:1966:EBE

[8] F. Stenger. Error bounds for the evaluation of integrals by repeated Gausstype formulae. Numerische Mathematik, 9(3):200–213, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1966:EBS

[9] Frank Stenger. Error Bounds for Solutions of Differential Equations. Ph.D. thesis, Department of Computing Science, University of Alberta, Edmonton, AB, Canada, 1966. 148 pp.

McNamee:1967:CFS

[10] J. McNamee and F. Stenger. Construction of fully symmetric numerical integration formulas. Numerische Mathematik, 10(4):327–344, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dolph:1968:SPH

[11] Charles L. Dolph, Frank Stenger, and A. Wiin-Nielsen. On the stability problems of the Helmholtz–Kelvin–Rayleigh type. Technical Report 08759-3-T, Department of Meteorology and Oceanography, University of Michigan, Ann Arbor, MI, USA, 1968. 72 pp.

Stenger:1968:BRB

[12] Frank Stenger. Book review: Numerical Integration (Philip J. Davis and Philip Rabinowitz). SIAM Review, 10(2):239–240, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1968:KPE

[13] Frank Stenger. Kronecker product extensions of linear operators. SIAM Journal on Numerical Analysis, 5(2):422–435, June 1968. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http:

//links.jstor.org/sici?sici=0036-1429(196806)5:2<422:KPEOLO>
2.0.CO%3B2-7.

Stenger:1968:RNI

[14] Frank Stenger. Review: Numerical Integration, by Philip J. Davis and Philip Rabinowitz. SIAM Review, 10(2):239–240, April 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/links.jstor.org/sici?sici=0036-1445(196804)10:2<239:NI>2.0.
CO%3B2-Y.

Stenger:1968:TC

[15] Frank Stenger. A trap in computations. ACM SIGNUM Newsletter, 3 (3):2, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [171].

Goodrich:1970:MSQ

[16] R. F. Goodrich and F. Stenger. Movable singularities and quadrature. Mathematics of Computation, 24(110):283–300, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/links.jstor.org/sici?sici=0025-5718(197004)24:110<283:MSAQ>
2.0.CO%3B2-S.

Stenger:1970:AAC

[17] Frank Stenger. The asymptotic approximation of certain integrals. SIAM Journal on Mathematical Analysis, 1(3):392–404, August 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stenger:1970:AST

[18] F. Stenger. On the asymptotic solution of two first order linear differential equations with large parameter. Funkcialaj Ekvacioj. Serio Internacia, 13: 1–18, 1970. CODEN FESIAT. ISSN 0532-8721. URL http://fe.math.

kobe-u.ac.jp/FE/Free/vol13/fe13-1.pdf.

G:1971:RTC

[19] W. G. Review: Tabulation of Certain Fully Symmetric Numerical Integration Formulas of Degree 7, 9 and 11, by Frank Stenger. Mathematics of Computation, 25(116):935, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2004361.

McNamee:1971:WCF

[20] J. McNamee, F. Stenger, and E. L. Whitney. Whittaker’s cardinal function in retrospect. Mathematics of Computation, 25(113):141–154, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://links.jstor.org/sici?sici=0025-
5718(197101)25:113<141:WCFIR>2.0.CO%3B2-6;   http://www.ams.
org/journals/mcom/1971-25-113/S0025-5718-1971-0301428-0/.

Stenger:1971:BRB

[21] Frank Stenger. Book review: Integrals and Sums (P. C. Chakravarti). SIAM Review, 13(4):582–583, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1971:CPA

[22] Frank Stenger. Constructive proofs for approximation by inner functions. Journal of Approximation Theory, 4(4):372–386, December 1971. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Stenger:1971:ERT

[23] Frank Stenger. Erratum: The reduction of two dimensional integrals into a finite number of one-dimensional integrals. Aequationes Mathematicae, 6(2–3):316–317, June 1971. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic). See [25].

Stenger:1971:RIS

[24] Frank Stenger. Review: Integrals and Sums, by P. C. Chakravarti. SIAM Review, 13(4):582–583, October 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.jstor.org/sici?

sici=0036-1445(197110)13:4<582:IAS>2.0.CO%3B2-1.

Stenger:1971:RTD

[25] Frank Stenger. The reduction of two dimensional integrals into a finite number of one dimensional integrals. Aequationes Mathematicae, 6(2–3): 278–287, June 1971. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic). See erratum [23].

Lipow:1972:HSC

[26] Peter R. Lipow and Frank Stenger. How slowly can quadrature formulas converge? Mathematics of Computation, 26(120):917–922, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://links.jstor.org/sici?sici=0025-5718(
197210)26:120<917:HSCQFC>2.0.CO%3B2-R.

Stenger:1972:ASW

[27] Frank Stenger. The approximate solution of Wiener–Hopf integral equations. Journal of Mathematical Analysis and Applications, 37(3):687–724, March 1972. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Stenger:1972:BRB

[28] Frank Stenger. Book review: Quadrature Formulae (A. Ghizetti and A. Ossicini). SIAM Review, 14(4):662, October 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1972:RQF

[29] Frank Stenger. Review: Quadrature Formulae, by A. Ghizetti and A. Ossicini. SIAM Review, 14(4):662, October 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.jstor.

org/sici?sici=0036-1445(197210)14:4<662:QF>2.0.CO%3B2-J.

Stenger:1972:TMO

[30] Frank Stenger. Transform methods for obtaining asymptotic expansions of definite integrals. SIAM Journal on Mathematical Analysis, 3(1):20–30, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stenger:1973:AAS

[31] Frank Stenger. An algorithm for the approximate solution of Wiener– Hopf integral equations. Communications of the ACM, 16(11):708–710, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stenger:1973:ASC

[32] Frank Stenger. The approximate solution of convolution-type integral equations. SIAM Journal on Mathematical Analysis, 4(3):536–555, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stenger:1973:BRB

[33] Frank Stenger. Book review: Approximate Calculation of Multiple Integrals (A. H. Stroud). SIAM Review, 15(1):234–235, January 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1973:IFB

[34] Frank Stenger. Integration formulae based on the trapezoidal formula. Journal of the Institute of Mathematics and its Applications, 12(1):103– 114, 1973. CODEN JMTAA8. ISSN 0020-2932. See remarks [53, 172].

Stenger:1973:RAC

[35] Frank Stenger. Review: Approximate Calculation of Multiple Integrals, by A. H. Stroud. SIAM Review, 15(1):234–235, January 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

links.jstor.org/sici?sici=0036-1445(197301)15:1<234:ACOMI>2.
0.CO%3B2-7;   https://epubs.siam.org/doi/abs/10.1137/1015023.

Gearhart:1974:ACE

[36] W. B. Gearhart and F. Stenger. An approximate convolution equation of a given response. In Kirby [198], pages 168–196. ISBN 3-540-07026-5. ISSN 0075-8442 (print), 2196-9957 (electronic). LCCN QA402.3 C33 1974.

Rahman:1974:EPP

[37] Q. I. Rahman and Frank Stenger. An extremal problem for polynomials with a prescribed zero. Proceedings of the American Mathematical Society, 43(1):84–90, March 1974. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://links.jstor.org/sici?

sici=0002-9939(197403)43:1<84:AEPFPW>2.0.CO%3B2-2.

Stenger:1974:CEB

[38] Frank Stenger. On the convergence and error of the Bubnov–Galerkin method. Lecture Notes in Mathematics, 362:434–450, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Stenger:1974:CTD

[39] Frank Stenger. Computing the topological degree of a mapping in R^{n}. Report, National Oceanic and Atmospheric Administration, Washington, DC, USA, 1974.

Chauvette:1975:ASN

[40] Jean Chauvette and Frank Stenger. The approximate solution of the nonlinear equation Δu = u − u^{3}. Journal of Mathematical Analysis and Applications, 51(1):229–242, July 1975. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Rosenberg:1975:LBA

[41] Ivo G. Rosenberg and Frank Stenger. A lower bound on the angles of triangles constructed by bisecting the longest side. Mathematics of Computation, 29(130):390–395, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.jstor.org/sici?

sici=0025-5718(197504)29:130<390:ALBOTA>2.0.CO%3B2-H.

Stenger:1975:AFW

[42] Frank Stenger. An analytic function which is an approximate characteristic function. SIAM Journal on Numerical Analysis, 12(2):239–254, April 1975. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170

(electronic). URL http://links.jstor.org/sici?sici=0036-1429(
197504)12:2<239:AAFWIA>2.0.CO%3B2-V.

Stenger:1975:ATD

[43] Frank Stenger. An algorithm for the topological degree of a mapping in nspace. Bulletin of the American Mathematical Society, 81(1):179–182, January 1975. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL http://projecteuclid.org/euclid.bams/1183536266.

Stenger:1975:CBC

[44] Frank Stenger. Connection between a Cauchy system representation of Kalaba and Fourier transforms. Applied Mathematics and Computation, 1(1):83–91, January 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stenger:1975:CTD

[45] Frank Stenger. Computing the topological degree of a mapping in R^{n}. Numerische Mathematik, 25(1):23–38, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1975:LBA

[46] F. Stenger and I. Rosenberg. A lower bound on the angles of triangles constructed by bisecting the longest side. Mathematics of Computation, 29 (130):390–395, 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.ams.org/journals/mcom/1975-29-
130/S0025-5718-1975-0375068-5/.

Harvey:1976:TDA

[47] Charles Harvey and Frank Stenger. A two-dimensional analogue to the method of bisections for solving nonlinear equations. Quarterly of Applied Mathematics, 33(??):351–368, ???? 1976. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic). URL http://www.ams.org/

journals/qam/1976-33-04/S0033-569X-1976-0455361-7/.

Ikebe:1976:NSH

[48] Y. Ikebe, T. Y. Li, and F. Stenger. The numerical solution of the Hilbert problem. In Theory of approximation, with applications (Proc. Conf., Univ. Calgary, Calgary, Alta., 1975; dedicated to the memory of Eckard Schmidt), pages 338–358. Academic Press Inc., New York, USA, 1976.

Nickel:1976:EBU

[49] Karl Nickel. Error bounds and uniqueness for the solutions of nonlinear, strongly coupled, parabolic systems of differential equations. MRC Technical Summary Report 1596, Mathematics Research Center, US Department of the Army, Madison, WI, USA, 1976. 20 pp.

Stenger:1976:ACE

[50] F. Stenger, W. Petrick, and Z. Rotsides. Algorithm for computing electromagnetic scattered field from an axially-symmetric body with an impedance boundary condition. SIAM Review, 18(4):828–829, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/doi/abs/10.1137/1018136.

Stenger:1976:AWC

[51] Frank Stenger. Approximations via Whittaker’s cardinal function. Journal of Approximation Theory, 17(3):222–240, July 1976. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Hanson:1977:DSS

[52] F. Hanson, J. M. Steele, and F. Stenger. Distinct sums over subsets. Proceedings of the American Mathematical Society, 66(1):179–180, September 1977. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

URL http://links.jstor.org/sici?sici=0002-9939(197709)66:1<
179:SNDSOS>2.0.CO%3B2-9;   http://www.ams.org/journals/proc/
1977-066-01/S0002-9939-1977-0447167-4/.

Stenger:1977:RIF

[53] Frank Stenger. Remarks on “Integration formulae based on the trapezoidal formula” (J. Inst. Math. Appl. 12 (1973), 103–114). Journal of the Institute of Mathematics and its Applications, 19(2):145–147, 1977. CODEN JMTAA8. ISSN 0020-2932. See [34, 172].

Stenger:1978:OCM

[54] Frank Stenger. Optimal convergence of minimum norm approximations in H_{p}. Numerische Mathematik, 29(4):345–362, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1978:ULE

[55] Frank Stenger. Upper and lower estimates on the rate of convergence of approximations in H_{p}. Bulletin of the American Mathematical Society, 84(1):145–148, 1978. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.org/

euclid.bams/1183540395.

Lundin:1979:CTA

[56] L. Lundin and F. Stenger. Cardinal-type approximations of a function and its derivatives. SIAM Journal on Mathematical Analysis, 10(1):139– 160, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-

7154 (electronic). URL https://epubs.siam.org/doi/abs/10.1137/
0510016.

Stenger:1979:SGM

[57] Frank Stenger. A “Sinc–Galerkin” method of solution of boundary value problems. Mathematics of Computation, 33(145):85–109, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://links.jstor.org/sici?sici=0025-5718(
197901)33:145<85:A%22MOSO>2.0.CO%3B2-S.

Stenger:1979:UTT

[58] Frank Stenger and Steven A. Johnson. Ultrasonic transmission tomography based on the inversion of the Helmholtz wave equation for plane and spherical wave insonification. Applied Mathematics Notes, 4(3–4):102–127, 1979. CODEN ???? ISSN 0700-9224.

Ball:1980:EIH

[59] James Ball, Steven A. Johnson, and Frank Stenger. Explicit inversion of the Helmholtz equation for ultra-sound insonification and spherical detection. In Wang [199], pages 451–461. ISBN 1-4684-3757-7, 1-4684-3755-0 (e-book). ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN QC1-75.

URL http://www.springerlink.com/content/978-1-4684-3755-3.

Stenger:1980:AES

[60] F. Stenger, M. Hagmann, and J. Schwing. An algorithm for the electromagnetic scattering due to an axially symmetric body with an impedance boundary condition. Journal of Mathematical Analysis and Applications, 78(2):531–573, December 1980. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic). URL https://www.sciencedirect.com/

science/article/pii/0022247X80901651.

Stenger:1981:AUT

[61] Frank Stenger. An algorithm for ultrasonic tomography based on inversion of the Helmholtz equation. In Allgower et al. [200], pages 371–406. CO-DEN LNMAA2. ISBN 0-387-10871-8, 3-540-10871-8 (print), 3-540-38781-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3

.L471 no.878; QA3.L471. URL http://link.springer.com/book/10.
1007/BFb0090674;   http://www.springerlink.com/content/978-3-
540-38781-7.

Stenger:1981:NMB

[62] Frank Stenger. Numerical methods based on Whittaker cardinal, or sinc functions. SIAM Review, 23(2):165–224, April 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.

jstor.org/sici?sici=0036-1445(198104)23:2<165:NMBOWC>2.0.CO%
3B2-S.

Johnson:1982:WEI

[63] Steven A. Johnson, Frank Stenger, Calvin Wilcox, James Ball, and Michael J. Berggren. Wave equations and inverse solutions for soft tissue. In Acoustical imaging, Vol. 11 (Monterey, Calif., 1981), pages 409–424. Plenum, New York, 1982.

Sikorski:1982:OQA

[64] K. Sikorski. Optimal quadrature algorithms in H_{p} spaces. Numerische Mathematik, 39(3):405–410, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/BF01407871.

Stenger:1982:AUI

[65] Frank Stenger. Asymptotic ultrasonic inversion based on using more than one frequency. In John P. Powers, editor, Acoustical imaging, Vol. 11 (Monterey, Calif., 1981), pages 425–444. Plenum, New York, NY, USA, 1982. ISBN 1-4684-1139-X, 1-4684-1137-3 (e-book). ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN QC1-75.

Stenger:1983:ANT

[66] Frank Stenger, Michael J. Berggren, Steven A. Johnson, and Y. Li. An adaptive, noise tolerant, frequency extrapolation algorithm for diffraction corrected ultrasound tomography. In McAvoy [201], pages 726–731. CODEN ULSPDT. ISBN ???? ISSN 0090-5607. LCCN ???? URL

http://ieeexplore.ieee.org/iel5/10283/32716/01535094.pdf?tp=
&arnumber=1535094&isnumber=32716. Two volumes. IEEE catalog num-

ber 83CH1947-1.

Eiger:1984:BMS

[67] A. Eiger, K. Sikorski, and F. Stenger. A bisection method for systems of nonlinear equations. ACM Transactions on Mathematical Software, 10(4): 367–377, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Johnson:1984:FIA

[68] S. A. Johnson, Y. Zhou, M. K. Tracy, M. J. Berggren, and F. Stenger. Fast iterative algorithms for inverse scattering solutions of the Helmholtz and Riccati wave equations. In Kaveh et al. [203], pages 75–87. CO-DEN ACIGD9. ISBN 0-306-41717-0. ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN ????

Johnson:1984:ISS

[69] S. A. Johnson, Y. Zhou, M. L. Tracey, M. J. Berggren, and F. Stenger. Inverse scattering solutions by a sinc basis, multiple source, moment method. Part III: Fast algorithms. Ultrasonic Imaging, 6(1):103–116, January 1984. CODEN ULIMD4. ISSN 0161-7346 (print), 1096-0910 (electronic).

Johnson:1984:RDS

[70] S. A. Johnson, M. J. Berggren, F. Stenger, C. H. Wilcox, and E. Jensen. Recent developments in solving the exact acoustic inverse scattering problem. In Anonymous, editor, Proceedings of the 29th Annual Meeting of the American Institute of Ultrasound in Medicine, and the 13th Annual Meeting of the Society of Diagnostic Medical Sonographers, 16–19 September 1984, Kansas City, MO, USA, pages 126–?? American Institute of Ultrasound in Medicine, Bethesda, MD, USA, 1984. ISBN ???? LCCN ????

Johnson:1984:UTG

[71] Steven A. Johnson and Frank Stenger. Ultrasound tomography by Galerkin or moment methods. In Nalcioğlu and Cho [196], pages 254–276. ISBN 3-642-93255-X, 3-642-93253-3 (e-book). ISSN 0172-7788. LCCN R858-859.7.

Sikorski:1984:AFS

[72] K. Sikorski, F. Stenger, and J. Schwing. Algorithm 614: A FORTRAN subroutine for numerical integration in H_{p}. ACM Transactions on Mathematical Software, 10(2):152–160, June 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sikorski:1984:FSI

[73] K. Sikorski, F. Stenger, and J. Schwing. A Fortran subroutine for integration in H_{p} spaces. ACM Transactions on Mathematical Software, 10(2):152–157, June 1984. CODEN ACMSCU. ISSN 0098-3500 (print),

1557-7295 (electronic). URL https://dl.acm.org/citation.cfm?doid=
399.449.

Sikorski:1984:OQS

[74] K. Sikorski and F. Stenger. Optimal quadratures in H_{p} spaces. ACM Transactions on Mathematical Software, 10(2):140–151, June 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stenger:1984:PSR

[75] Frank Stenger. Polynomial, sinc and rational function methods for approximating analytic functions. Lecture Notes in Mathematics, 1105:49– 72, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072399/.

Stenger:1984:RFF

[76] F. Stenger, M. J. Berggren, S. A. Johnson, and C. H. Wilcox. Rational function frequency extrapolation in ultrasonic tomography. In Wave phenomena: modern theory and applications (Toronto, 1983), volume 97 of North-Holland Math. Stud., pages 19–34. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984.

Stenger:1984:SMA

[77] Frank Stenger. Sinc methods of approximate solution of partial differential equations. In Miller [204], pages 40–64. ISBN 0-86784-508-2. LCCN QA374 .C661 1984.

Kim:1985:ISS

[78] W. W. Kim, Michael J. Berggren, Steven A. Johnson, Frank Stenger, and Calvin H. Wilcox. Inverse scattering solutions to the exact Riccati wave equations by iterative RYTOV approximations and internal field calculations. In McAvoy [205], pages 878–882. CODEN UL-SPDT. ISBN ???? ISSN 0090-5607. LCCN TA367 U47 1985 v. 1-

2. URL http://ieeexplore.ieee.org/iel5/10285/32718/01535578.
pdf?tp=&arnumber=1535578&isnumber=32718. Two volumes. IEEE cat-

alog number 85CH2209-5.

Berggren:1986:UIS

[79] M. J. Berggren, S. A. Johnson, B. L. Carruth, W. W. Kim, F. Stenger, and P. K. Kuhn. Ultrasound inverse scattering solutions from transmission and/or reflection data. In Nalcioglu et al. [206], pages 114–121. CODEN PSISDG. ISBN 0-89252-706-4. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK8315 .I5451 1986; TK8315 .I57 1986; TA1632 .I62 1986; TS510 .S63; TA1632 .I58 1986.

Schaffer:1986:MSM

[80] Steve Schaffer and Frank Stenger. Multigrid-sinc methods. Applied Mathematics and Computation, 19(1–4):311–319, July 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Stenger:1986:ENO

[81] Frank Stenger. Explicit, nearly optimal, linear rational approximation with preassigned poles. Mathematics of Computation, 47(175):225–252, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://links.jstor.org/sici?sici=0025-5718(
198607)47:175<225:ENOLRA>2.0.CO%3B2-P.

Berggren:1987:PFI

[82] M. J. Berggren, S. A. Johnson, B. L. Carruth, W. W. Kim, F. Stenger, and P. L. Kuhn. Performance of fast inverse scattering solutions for the exact Helmholtz equation using multiple frequencies and limited views. In Jones [207], pages 193–201. CODEN ACIGD9. ISBN 0-306-42565-3. ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN ????

Kearfott:1987:SFF

[83] R. B. Kearfott, K. Sikorski, and F. Stenger. A Sinc function fast Poisson solver, 1987.

Kim:1987:AIS

[84] W. W. Kim, S. A. Johnson, M. J. Berggren, F. Stenger, and C. H. Wilcox. Analysis of inverse scattering solutions from single frequency, combined transmission and reflection data for the Helmholtz and Riccati exact wave equations. In Jones [207], pages 359–369. CODEN ACIGD9. ISBN 0-306-42565-3. ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN ????

Ang:1988:NTP

[85] – D.ǎng – Diñh Áng, Fritz Keinert, and Frank Stenger. A nonlinear two-phase Stefan problem with melting point gradient: a constructive approach. Journal of Computational and Applied Mathematics, 23(2):245–255, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
037704278890283X.

Bialecki:1988:SNM

[86] Bernard Bialecki and Frank Stenger. Sinc–Nyström method for numerical solution of one-dimensional Cauchy singular integral equation given on a smooth arc in the complex plane. Mathematics of Computation, 51(183):133–165, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.jstor.org/sici?

sici=0025-5718(198807)51:183<133:SMFNSO>2.0.CO%3B2-7.

Dikshit:1988:RTW

[87] H. P. Dikshit, A. Sharma, V. Singh, and F. Stenger. Rivlin’s theorem on Walsh equiconvergence. Journal of Approximation Theory, 52(3):339– 349, March 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Stenger:1988:BRB

[88] Frank Stenger. Book review: Computational Complexity (K. Wagner and G. Wechsung). SIAM Review, 30(2):353–354, June 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/links.jstor.org/sici?sici=0036-1445(198806)30:2<353:CC>2.0.
CO%3B2-B.

Stenger:1988:RCC

[89] Frank Stenger. Review: Computational Complexity, by K. Wagner and G. Wechsung. SIAM Review, 30(2):353–354, June 1988. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.

jstor.org/sici?sici=0036-1445(198806)30:2<353:CC>2.0.CO%3B2-
B;   https://epubs.siam.org/doi/abs/10.1137/1030086.

Ang:1989:CVR

[90] – D.ǎng – Diñh Áng, John Lund, and Frank Stenger. Complex variable and regularization methods of inversion of the Laplace transform. Mathematics of Computation, 53(188):589–608, October 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.

jstor.org/sici?sici=0025-5718(198910)53:188<589:CVARMO>2.0.CO%
3B2-H.

Ang:1989:VFD

[91] – D.ǎng – Diñh Áng, Tim Folias, Fritz Keinert, and Frank Stenger. Viscoplastic flow due to penetration: a free boundary value problem. International Journal of Fracture, 39(1–3):121–127, 1989. CODEN IJFRAP. ISSN 0376-9429 (print), 1573-2673 (electronic).

Kowalski:1989:OCR

[92] Marek A. Kowalski and Frank Stenger. Optimal complexity recovery of band- and energy-limited signals. II. Journal of Complexity, 5(1):45–59, March 1989. CODEN JOCOEH. ISSN 0885-064X (print), 1090-2708 (electronic).

Stenger:1989:EAM

[93] Frank Stenger. Explicit approximate methods for computational control theory. In Bowers et al. [208], pages 299–316. ISBN 0-8176-3438-X. LCCN TA329 .C645 1989.

Ikebe:1990:RAS

[94] Yasuhiko Ikebe, Marek Kowalski, and Frank Stenger. Rational approximation of the step, filter, and impulse functions. In Wong [210], pages 441–454. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL http:/

/www.loc.gov/catdir/enhancements/fy0647/90002810-d.html.

Stenger:1990:BRB

[95] Frank Stenger. Book review: Rational Approximation of Real Functions (P. P. Petrushev and V. I. Popov). SIAM Review, 32(1):187–188, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200

(electronic). URL http://links.jstor.org/sici?sici=0036-1445(
199003)32:1<187:RAORF>2.0.CO%3B2-R.

Stenger:1990:RRA

[96] Frank Stenger. Review: Rational Approximation of Real Functions by P. P. Petrushev and V. I. Popov. SIAM Review, 32(1):187–188, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://links.jstor.org/sici?sici=0036-1445(199003)32:1<
187:RAORF>2.0.CO%3B2-R;   https://epubs.siam.org/doi/abs/10.
1137/1032034.

Stenger:1990:SOR

[97] Frank Stenger. Some open research problems in sonic and electromagnetic inversion. In Martin and White [209], pages 73–89.

Gustafson:1991:CAA

[98] Sven-Åke Gustafson and Frank Stenger. Convergence acceleration applied to Sinc approximation with application to approximation of |x|^{α}. In Bowers and Lund [211], pages 161–171. ISBN 0-8176-3611-0. LCCN TA329 .C644 1991. US$65.00.

Stenger:1992:SII

[99] Frank Stenger, Brian Keyes, Mike O’Reilly, and Ken Parker. Sinc indefinite integration and initial value problems. In Espelid and Genz [212], pages 281–282. ISBN 0-7923-1583-9. LCCN QA299.3 .N38 1991.

Stenger:1993:BRB

[100] Frank Stenger. Book review: Sinc Methods for Quadrature and Differential Equations (J. Lund and K. L. Bowers). SIAM Review, 35(4):682– 683, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://links.jstor.org/sici?sici=0036-
1445(199312)35:4<682:SMFQAD>2.0.CO%3B2-F.

Stenger:1993:DE

[101] Frank Stenger. Differential equations. In Numerical methods based on Sinc and analytic functions [213], chapter 7, pages 441–532. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Stenger:1993:IE

[102] Frank Stenger. Integral equations. In Numerical methods based on Sinc and analytic functions [213], chapter 6, pages 311–440. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Stenger:1993:MP

[103] Frank Stenger. Mathematical preliminaries. In Numerical methods based on Sinc and analytic functions [213], chapter 1, pages 1–103. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Stenger:1993:PA

[104] Frank Stenger. Polynomial approximation. In Numerical methods based on Sinc and analytic functions [213], chapter 2, pages 105–130. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Stenger:1993:RSM

[105] Frank Stenger. Review: Sinc Methods for Quadrature and Differential Equations, by J. Lund and K. L. Bowers. SIAM Review, 35 (4):682–683, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.jstor.org/sici?

sici=0036-1445(199312)35:4<682:SMFQAD>2.0.CO%3B2-F;   https:
//epubs.siam.org/doi/abs/10.1137/1035172.

Stenger:1993:SA

[106] Frank Stenger. Sinc approximation on Γ. In Numerical methods based on Sinc and analytic functions [213], chapter 4, pages 179–242. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Stenger:1993:SAS

[107] Frank Stenger. Sinc approximation in strip. In Numerical methods based on Sinc and analytic functions [213], chapter 3, pages 131–178. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Stenger:1993:SCA

[108] F. Stenger, B. Barkey, and R. Vakili. Sinc convolution approximate solution of Burgers’ equation. In Computation and control, III (Bozeman, MT, 1992), volume 15 of Progr. Systems Control Theory, pages 341–354. Birkhäuser Boston Inc., Cambridge, MA, USA, 1993.

Stenger:1993:SRM

[109] Frank Stenger. Sinc-related methods. In Numerical methods based on Sinc and analytic functions [213], chapter 5, pages 243–310. ISBN 978-146-127-6-3-7-1. LCCN QA372 .S82 1993.

Chan:1994:NPB

[110] Kwong-Yu Chan, Douglas Henderson, and Frank Stenger. Nonlinear Poisson–Boltzmann equation in a model of a scanning tunneling microscope. Numerical Methods for Partial Differential Equations, 10(6):689– 702, 1994. CODEN NMPDEB. ISSN 0749-159X (print), 1098-2426 (electronic).

McArthur:1994:RNM

[111] K. M. McArthur. Review: Numerical Methods Based on Sinc and Analytic Functions (Frank Stenger). SIAM Review, 36(4): 673–674, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.jstor.org/sici?

sici=0036-1445(199412)36:4<673:NMBOSA>2.0.CO%3B2-J;   https:
//epubs.siam.org/doi/abs/10.1137/1036167.

Schmeisser:1994:RBM

[112] G. Schmeisser. Review: Numerical Methods Based on Sinc and Analytic Functions, by Frank Stenger. Mathematics of Computation, 63 (208):817–819, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.jstor.org/sici?

sici=0025-5718(199410)63:208<817:NMBOSA>2.0.CO%3B2-K.

Stenger:1994:NMT

[113] Frank Stenger. Numerical methods via transformations. In Zahar [214], pages 543–550. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Kowalski:1995:STA

[114] Marek A. Kowalski, Krzysztof A. Sikorski, and Frank Stenger. Selected Topics in Approximation and Computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-508059-9. xiv + 349

pp. URL http://site.ebrary.com/lib/utah/Doc?id=10087215.

Morlet:1995:SAS

[115] Anne C. Morlet and Frank Stenger. Sinc approximation of solution of heat equation with discontinuous initial condition. In Computation and control, IV (Bozeman, MT, 1994), volume 20 of Progr. Systems Control Theory, pages 289–303. Birkhäuser Boston Inc., Cambridge, MA, USA, 1995.

Stenger:1995:CC

[116] Frank Stenger. Collocating convolutions. Mathematics of Computation, 64(209):211–235, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.jstor.org/sici?

sici=0025-5718(199501)64:209<211:CC>2.0.CO%3B2-M.

Stenger:1995:SCT

[117] Frank Stenger. Sinc convolution — a tool for circumventing some limitations of classical signal processing. In Ismail et al. [216], pages 227–

240. ISBN 0-8218-0384-0. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA299.6 .M38 1995.

Stenger:1995:SIH

[118] Frank Stenger. Sinc inversion of the Helmholtz equation without computing the forward solution. In Áng et al. [215], pages 149–157. ISBN ???? LCCN ????

Stenger:1996:BRS

[119] Frank Stenger. Book reviews: Solving Problems in Scientific Computing Using MAPLE and MATLAB, by Walter Gander and Jírí Hrebícek. Mathematics of Computation, 65(214):880–882, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:

//www.ams.org/jourcgi/jour-pbprocess?fn=110&arg1=S0025-5718-
96-00700-4&u=/mcom/1996-65-214/.

Bojanov:1997:BRS

[120] Borislav Bojanov. Book review: Selected Topics in Approximation and Computation, by Marek A. Kowalski, Krzysztof A. Sikorski, and Frank Stenger. SIAM Review, 39(2):333–334, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://links.

jstor.org/sici?sici=0036-1445(199706)39:2<333:STIAAC>2.0.CO%
3B2-Y;   https://epubs.siam.org/doi/abs/10.1137/SIREAD000039000002000333000001.

Quak:1997:BRS

[121] Ewald Quak. Book reviews: Selected topics in approximation and computation, by Marek A. Kowalski, Krzysztof A. Sikorski and Frank Stenger. Mathematics of Computation, 66(219):1374, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://links.jstor.org/sici?sici=0025-5718(
199707)66:219<1374:STIAAC>2.0.CO%3B2-L;   http://www.ams.org/
journals/mcom/1997-66-219/S0025-5718-97-00877-6/.

Stenger:1997:MSM

[122] Frank Stenger. Matrices of Sinc methods. Journal of Computational and Applied Mathematics, 86(1):297–310, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0377042797001635.

Special issue dedicated to William B. Gragg (Monterey, CA, 1996).

Stenger:1997:RDT

[123] Frank Stenger. Reviews and descriptions of tables and books: 22. Integral equations: Theory and numerical treatment, by Wolfgang Hackbusch. Mathematics of Computation, 66(220):1756–1758, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/jourcgi/jour-pbprocess?fn=110&
arg1=S0025-5718-97-00843-0&u=/mcom/1997-66-220/.

Stenger:1997:RIE

[124] Frank Stenger. Review: Integral Equations: Theory and Numerical Treatment, by Wolfgang Hackbusch. Mathematics of Computation, 66(220):1756–1758, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.

jstor.org/sici?sici=0025-5718(199710)66:220<1756:IETANT>2.0.
CO%3B2-V;   http://www.ams.org/journals/mcom/1997-66-220/S0025-
5718-97-00910-1/;   https://www.jstor.org/stable/2153702.

Chen:1998:HSS

[125] Susheela Narasimhan Kuan Chen and Frank Stenger. A harmonic-sinc solution of the Laplace equation for problems with singularities and semiinfinite domains. Numerical Heat Transfer, Part B (Fundamentals), 33 (4):433–450, June 1998. CODEN NUHTD6. ISSN 1040-7790 (print), 1521-0626 (electronic).

Kress:1998:SQM

[126] Rainer Kress, Ian H. Sloan, and Frank Stenger. A sinc quadrature method for the double-layer integral equation in planar domains with corners. Journal of Integral Equations and Applications, 10(3):291–317, 1998. CO-DEN ???? ISSN 0897-3962 (print), 1938-2626 (electronic). URL http:/

/projecteuclid.org/euclid.jiea/1181074232.

Narasimhan:1998:HSS

[127] S. Narasimhan, Kuan Chen, and Frank Stenger. A harmonic-sinc solution of the Laplace equation for problems with singularities and semiinfinite domains. Numerical Heat Transfer, Part B (Fundamentals), 33 (4):433–450, June 1998. CODEN NUHTD6. ISSN 1040-7790 (print),

1521-0626 (electronic). URL https://www.tandfonline.com/doi/abs/
10.1080/10407799808915042.

Stenger:1998:CSM

[128] Frank Stenger and Michael J. O’Reilly. Computing solutions to medical problems via Sinc convolution. IEEE Transactions on Automatic Control, 43(6):843–848, June 1998. CODEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Stenger:1999:BRB

[129] Frank Stenger. Book reviews: Boundary element method, fundamentals and applications, by Frederico Paris and Jose Canas. Mathematics of Computation, 68(225):457–459, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.jstor.org/

sici?sici=0025-5718(199901)68:225<457:BEMFAA>2.0.CO%3B2-K;
http://www.ams.org/jourcgi/jour-pbprocess?fn=110&arg1=S0025-
5718-99-00992-8&u=/mcom/1999-68-225/.

Stenger:1999:CMS

[130] Frank Stenger and Ross Schmidtlein. Conformal maps via Sinc methods. In Papamichael et al. [217], pages 505–549. ISBN 981-02-3626-3. LCCN QA297 .I473 1997.

Stenger:1999:OIP

[131] F. Stenger, S.-Å. Gustafson, B. Keyes, M. O’Reilly, and K. Parker. ODE-IVP-PACK via Sinc indefinite integration and Newton’s method. Numerical Algorithms, 20(2–3):241–268, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/18/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/18/7/fulltext.
pdf;   https://link.springer.com/article/10.1023/A%3A1019108002140.

Narasimhan:2000:SIN

[132] S. Narasimhan, Kuan Chen, and F. Stenger. The solution of incompressible Navier–Stokes equations using the sine collocation method. In Kromann et al. [218], pages 199–214. ISBN 0-7803-5912-7 (softcover), 0-7803-5913-5 (casebound), 0-7803-5914-3 (microfiche). LCCN TK7870.25 .I6 2000. Two volumes. IEEE catalog number 00CH37069.

Resch:2000:FER

[133] Ron Resch, Frank Stenger, and Jörg Waldvogel. Functional equations related to the iteration of functions. Aequationes Mathematicae, 60(1– 2):25–37, August 2000. CODEN AEMABN. ISSN 0001-9054 (print),

1420-8903 (electronic). URL https://link.springer.com/article/10.
1007/s000100050133.

Stenger:2000:NSS

[134] F. Stenger, R. Chaudhuri, and J. Chiu. Novel sinc solution of the boundary integral form for two-dimensional bi-material elasticity problems. Composites Science and Technology, 60(12–13):2197–2211, September 2000. CODEN CSTCEH. ISSN 0266-3538 (print), 1879-1050 (electronic).

Stenger:2000:SAC

[135] Frank Stenger. Sinc approximation of Cauchy-type integrals over arcs. The ANZIAM Journal, 42(1):87–97, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/sinc-approximation-
of-cauchytype-integrals-over-arcs/F3E8E58B66C08F142BB6C63D61317703.

Papers in honour of David Elliott on the occasion of his sixty-fifth birthday.

Stenger:2000:SSN

[136] Frank Stenger. Summary of Sinc numerical methods. Journal of Computational and Applied Mathematics, 121(1–2):379–420, September 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/

S0377042700003484. Numerical analysis in the 20th century, Vol. I, Approximation theory.

Stenger:2000:TDH

[137] Frank Stenger. Three dimensional hybrid BEM–Sinc analysis of bonded/ bolted composite joints with discrete cracks. Technical Report AD-a376 152, SIN-0005, Sinc. Inc., Salt Lake City, UT, USA, 2000. 53 pp.

Stenger:2000:UAA

[138] Frank Stenger. A unified approach to the approximate solution of PDE. Berichte aus der Technomathematik 00-17, Zentrum für Technomathematik, Bremen, Germany, 2000. 47 pp.

Narasimhan:2002:FSA

[139] S. Narasimhan, Kuan Chen, and Frank Stenger. A first step in applying the Sinc collocation method to the nonlinear Navier–Stokes equations. Numerical Heat Transfer, Part B (Fundamentals), 41(5):447–462, May 1, 2002. CODEN NUHTD6. ISSN 1040-7790 (print), 1521-0626 (electronic).

Stenger:2002:RFM

[140] Frank Stenger, Elaine Cohen, and Richard Riesenfeld. Radial function methods of approximation based on using harmonic Green’s functions. Communications in Applied Analysis, 6(1):1–15, 2002. CODEN ???? ISSN 1083-2564.

Stenger:2002:SMA

[141] Frank Stenger, Ahmad Reza Naghsh-Nilchi, Jenny Niebsch, and Ronny Ramlau. Sampling methods for approximate solution of PDE. In Nashed and Scherzer [219], pages 199–249. ISBN 0-8218-2979-3. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN TA1637 .A47 2001.

Chambers:2003:ERB

[142] Janice J. Chambers, Shaheed Almudhafar, and Frank Stenger. Effect of reduced beam section frame elements on stiffness of moment frames. Journal of Structural Engineering, 129(3):383–393, March 2003. CODEN JSENEI. ISSN 0970-0137.

Stenger:2004:SCS

[143] Frank Stenger. Sinc convolution solution of laminated and anisotropic composite joints with discrete cracks. Technical Report AD-b302 193, Sinc. Inc., Salt Lake City, UT, USA, 2004. 473 pp.

Stenger:2004:SSB

[144] Frank Stenger, Thomas Cook, and Robert M. Kirby. Sinc solution of biharmonic problems. Canadian Applied Mathematics Quarterly, 12(3): 391–414, Fall 2004. CODEN ???? ISSN 1073-1849 (print), 1938-2634

(electronic). URL http://www.math.ualberta.ca/ami/CAMQ/table_
of_content/vol_12/12_3h.htm.

Schmeisser:2007:SAG

[145] Gerhard Schmeisser and Frank Stenger. Sinc approximation with a Gaussian multiplier. Sampling Theory in Signal and Image Processing, 6(2):199–221, 2007. ISSN 1530-6429. URL http://stsip.org/pdf_

campaign/vol06/no2/vol06no2pp199-221.pdf.

Stenger:2007:SVS

[146] Frank Stenger. Separation of variables solution of PDE via sinc methods. International Journal of Applied Mathematics & Statistics, 10(SO7):98– 115, 2007. ISSN 0973-1377 (print), 0973-7545 (electronic). URL http://

www.ceser.in/ceserp/index.php/ijamas/article/view/305.

Stenger:2008:FSZ

[147] Frank Stenger. Fourier series for zeta function via Sinc. Linear Algebra and its Applications, 429(10):2636–2639, November 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stenger:2008:PAB

[148] Frank Stenger, Brandon Baker, Carl Brewer, Geoffrey Hunter, Sasha Kaputerko, and Jason Shepherd. Periodic approximations based on sinc. International Journal of Pure and Applied Mathematics, 49(1):63–72, 2008. ISSN 1311-8080 (print), 1314-3395 (electronic). URL http://ijpam.eu/

contents/2008-49-1/8/index.html.

Stenger:2009:PFD

[149] Frank Stenger. Polynomial function and derivative approximation of Sinc data. Journal of Complexity, 25(3):292–302, 2009. CODEN JOCOEH. ISSN 0885-064X (print), 1090-2708 (electronic).

Baumann:2011:FCS

[150] Gerd Baumann and Frank Stenger. Fractional calculus and Sinc methods. Fractional Calculus and Applied Analysis, 14(4):568–622, 2011. ISSN 1311-0454 (print), 1314-2444 (electronic).

Stenger:2011:HSN

[151] Frank Stenger. Handbook of Sinc Numerical Methods. Chapman and Hall/CRC numerical analysis and scientific computation series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-2158-5 (hardback), 1-4398-2159-3 (e-book). xx + 463 pp. LCCN QA372 .S8195 2010. URL http://www.crcpress.com/

product/isbn/9781439821589.

Baumann:2013:FAD

[152] Gerd Baumann and Frank Stenger. Fractional adsorption diffusion. Fractional Calculus and Applied Analysis, 16(3):737–764, 2013. ISSN 1311-0454 (print), 1314-2444 (electronic).

Hagmann:2013:LHM

[153] Mark J. Hagmann, Frank S. Stenger, and Dmitry A. Yarotski. Linewidth of the harmonics in a microwave frequency comb generated by focusing a mode-locked ultrafast laser on a tunneling junction. Journal of Applied Physics, 114(22):223107, 2013. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://scitation.aip.

org/content/aip/journal/jap/114/22/10.1063/1.4831952.

Stenger:2013:IAU

[154] Frank Stenger, Maha Youssef, and Jenny Niebsch. Improved approximation via use of transformations. In Shen and Zayed [220], chapter 2, pages 25–49. ISBN 1-4614-4144-7, 1-4614-4145-5 (e-book). LCCN TA342 .M855 2013.

Stenger:2013:SHR

[155] Frank Stenger. Some history and research of Frank Stenger. Web doc-

ument, April 10, 2013. URL https://users.cs.utah.edu/~stenger/
history.pdf.

Stenger:2013:SMC

[156] Frank Stenger and Richard B. Hall. Sinc methods for computing solutions to viscoelastic and related problems. Canadian Applied Mathematics Quarterly, 21(1):95–120, 2013. ISSN 1073-1849 (print), 1938-2634 (electronic).

Stenger:2014:LCS

[157] Frank Stenger, Hany A. M. El-Sharkawy, and Gerd Baumann. The Lebesgue constant for sinc approximations. In Zayed and Schmeisser [221], pages 319–335. ISBN 3-319-08800-9, 3-319-08801-7 (e-book). ISSN 2296-5009 (print), 2296-5017 (electronic). LCCN ????

Baumann:2015:SAF

[158] Gerd Baumann and Frank Stenger. Sinc-approximations of fractional operators: a computing approach. Mathematics, 3(2):444–480, 2015. ISSN 2227-7390.

Stenger:2015:CMCa

[159] Frank Stenger, Gerd Baumann, and Vasilios G. Koures. Computational methods for chemistry and physics, and schrödinger in 3 + 1. Preprint, University of Utah; German University in Cairo; IISAM L3C, Salt Lake City, UT, USA; New Cairo City, Egypt; Oklahoma City, OK, USA, February 24, 2015. 44 pp. To be published in the proceedings of a conference of December 2014 honoring the 85th birthday of Frank E. Harris.

Stenger:2015:CMCb

[160] Frank Stenger, Gerd Baumann, and Vasilios G. Koures. Computational methods for chemistry and physics, and Schrödinger in 3+1^{1}. Advances in Quantum Chemistry, 71:265–298, 2015. CODEN AQCHA9. ISSN 0065-3276 (print), 2162-8815 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0065327615000064.

Stenger:2016:NSE

[161] Frank Stenger, Don Tucker, and Gerd Baumann. Navier–Stokes Equations on R^{3} ×[0, T]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-27524-0 (hardcover), 3-319-27526-7 (e-book). xi + 219 pp. LCCN QA374 .S834 2016. URL http://www.

springer.com/us/book/9783319275246.

Baumann:2017:FFP

[162] Gerd Baumann and Frank Stenger. Fractional Fokker–Planck equation. Mathematics, 5(1):19, 2017. ISSN 2227-7390.

Stenger:2017:ARZ

[163] Frank Stenger. Approximating the Riemann zeta and related functions. In Govil et al. [224], pages 363–373. ISBN 3-319-49240-3 (print), 3-319-49242-X (e-book). LCCN QA402.5-402.6. URL https://link.springer.

com/chapter/10.1007/978-3-319-49242-1_17.

Stenger:2018:ASN

[164] Frank Stenger. Approximate solution of numerical problems via approximate indefinite integration. Talk dedicated to Walter Gautschi on the occasion of his 90th birthday., April 4, 2018. PDF file not yet released.

Stenger:2018:IIO

[165] Frank Stenger. Indefinite integration operators identities and their approximations. arXiv.org, pages 1–36, October 17, 2018. URL https://

arxiv.org/abs/1809.05607.

Stenger:2018:PRH

[166] Frank Stenger. A proof of the Riemann hypothesis. arXiv.org, pages 1–26,

February 8, 2018. URL https://arxiv.org/abs/1708.01209v4.

Stenger:2021:IIO

[167] Frank Stenger. Indefinite integration operators identities, and their approximations. In Baumann [225], chapter 9, pages 227–254. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Stenger:2022:EPR

[168] Frank Stenger. An elementary proof of the Riemann Hypothesis. Report, School of Computing, University of Utah, Salt Lake City, UT 84112, USA,

April 6, 2022. 13 pp. URL https://arxiv.org/abs/1708.01209.

Stenger:2023:AZR

[169] Frank Stenger. All zeros of the Riemann zeta function in the critical strip are located on the critical line and are simple. Advances in Pure and Applied Mathematics, 13(6):204–411, June 29, 2023. ISSN 1867-1152 (print), 1869-6090 (electronic). URL https://www.scirp.org/pdf/apm_

2023062814170870.pdf.

Stenger:2024:AZR

[170] Frank Stenger. All zeros of the Riemann zeta function in the critical strip are located on the critical line and are simple. arXiv.org, ??(??):1–16, July

29, 2024. URL https://arxiv.org/abs/1708.01209.

Moler:1969:MSC

[171] C. B. Moler. More on the sphere in the corner. ACM SIGNUM Newsletter, 4(1):7, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [15].

Sack:1978:CSQ

[172] R. A. Sack. Comments on some quadrature formulas by F. Stenger. Journal of the Institute of Mathematics and its Applications, 21(3):359– 361, April 1978. CODEN JMTAA8. ISSN 0020-2932. URL https://

academic.oup.com/imamat/article/21/3/359/690581. See [34, 53].

Stynes:1979:SST

[173] Martin Stynes. A simplification of Stenger’s topological degree formula. Numerische Mathematik, 33(2):147–155, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/BF01399550. See [45].

Beighton:1982:EES

[174] S. Beighton and B. Noble. An error estimate for Stenger’s quadrature formula. Mathematics of Computation, 38(158):539–545, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/mcom/1982-38-158/S0025-5718-
1982-0645669-9/. See [34].

Stynes:1982:RST

[175] M. Stynes. A remark on Stenger’s topological degree algorithm. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 82(2):163–166, 1982. ISSN 0035-8975. URL https://www.

jstor.org/stable/20489150.

McArthur:1994:BRB

[176] Kelly M. McArthur. Book review: Numerical Methods Based on Sine and Analytic Functions, by Frank Stenger. SIAM Review, 36(4):673–674, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [213].

Davis:2004:IFS

[177] Philip Davis. An interview with Frank Stenger. Interview conducted by the Society for Industrial and Applied Mathematics, as part of grant # DE-FG02-01ER25547 awarded by the US Department of Energy, June 24,

2004. URL http://history.siam.org/oralhistories/stenger.htm.

Bialecki:2009:GEP

[178] Bernard Bialecki, Baker R. Kearfott, Krzysztof A. Sikorski, and Masaki Sugihara. Guest Editors’ preface: Issue dedicated to Professor Frank Stenger. Journal of Complexity, 25(3):233–236, June 2009. CODEN JO-COEH. ISSN 0885-064X (print), 1090-2708 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0885064X09000053.

Han:2014:PSC

[179] Lixing Han and Jianhong Xu. Proof of Stenger’s conjecture on matrix I^{(−1)} of Sinc methods. Journal of Computational and Applied Mathematics, 255(??):805–811, January 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0377042713003452. See [122].

Gautschi:2019:CST

[180] Walter Gautschi and Ernst Hairer. On conjectures of Stenger in the theory of orthogonal polynomials. Journal of Inequalities and Applications, page 27, 2019. ISSN 1025-5834. URL https://

journalofinequalitiesandapplications.springeropen.com/articles/
10.1186/s13660-019-2107-6. Paper no. 159.

Annaby:2021:OCE

[181] M. H. Annaby, R. M. Asharabi, and M. M. Tharwat. An overview of the computation of the eigenvalues using sinc-methods. In Baumann [225], chapter 10, pages 255–298. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Annaby:2021:SGA

[182] M. H. Annaby and R. M. Asharabi. Sinc-Gaussian approach for solving the inverse heat conduction problem. In Baumann [225], chapter 1, pages 3–21. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Attili:2021:NSF

[183] Basem Attili. Numerical solution of the Falkner–Skan equation arising in boundary layer theory using the sinc-collocation method. In Baumann [225], chapter 7, pages 147–162. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Baumann:2021:LSE

[184] Gerd Baumann. Lévy–Schrödinger equation: Their eigenvalues and eigenfunctions using sinc methods. In New Sinc methods of numerical analysis: Festschrift in honor of Frank Stenger’s 80th birthday [225], chapter 4, pages 55–98. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Beebe:2021:PAF

[185] Nelson H. F. Beebe. Publications by, and about, Frank Stenger. In Baumann [225], pages 385–399. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Berrut:2021:IJS

[186] Jean-Paul Berrut. The influence of jumps on the sinc interpolant, and ways to overcome it. In Baumann [225], chapter 12, pages 323–339. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Dopp:2021:EIA

[187] Kathy Dopp. Election integrity audits to ensure election outcome accuracy. In Baumann [225], chapter 6, pages 123–145. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Ismail:2021:CMF

[188] Mourad E. H. Ismail and Ruiming Zhang. Completely monotonic Fredholm determinants. In Baumann [225], chapter 11, pages 299–321. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Nedaiasl:2021:SPS

[189] Khadijeh Nedaiasl. Sinc projection solutions of Fredholm integral equations. In Baumann [225], chapter 3, pages 35–53. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Okayama:2021:MSQ

[190] Tomoaki Okayama and Shu Hanada. A modified Stenger’s quadrature formula for infinite integrals of unilateral rapidly decreasing functions and its theoretical error bound. Mathematics and Computers in Simulation, 186(??):3–18, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.sciencedirect.com/

science/article/pii/S037847542030080X.

Rashidinia:2021:ASM

[191] J. Rashidinia, A. Parsa, and R. Salehi. Application of sinc on the multiorder fractional differential equations. In Baumann [225], chapter 5, pages 99–122. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Stromberg:2021:FM

[192] Marc Stromberg. LU factorization of any matrix. In Baumann [225], chapter 14, pages 369–382. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Stromberg:2021:SMP

[193] Marc Stromberg. Sinc methods on polyhedra. In Baumann [225], chapter 8, pages 163–224. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Tanaka:2021:CAF

[194] Ken’ichiro Tanaka and Masaaki Sugihara. Construction of approximation formulas for analytic functions by mathematical optimization. In Baumann [225], chapter 13, pages 341–368. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Youssef:2021:PSC

[195] Maha Youssef. Poly-sinc collocation method for solving coupled Burgers equations with a large Reynolds number. In Baumann [225], chapter 2, pages 23–34. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Nalcioglu:1984:STI

[196] Orhan Nalcioğlu and Zang-Hee Cho, editors. Selected Topics in Image Science, volume 23 of Lecture Notes in Medical Informatics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-642-93255-X, 3-642-93253-3 (e-book). ISSN 0172-7788. ???? pp. LCCN R858-859.7.

Bettis:1974:PCN

[197] Dale G. Bettis, editor. Proceedings of the Conference on the Numerical Solution of Ordinary Differential Equations: 19,20 October 1972, The University of Texas at Austin, volume 362 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 0-387-06602-0, 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L35; QA3 .L28 no. 362; QA372; QA3 .L28; QA1 .L471; QA3 .L4; QA372 .C765p

1972. URL http://link.springer.com/book/10.1007/BFb0066582;
http://www.springerlink.com/content/978-3-540-37911-9.

Kirby:1974:OCT

[198] Bruce J. Kirby, editor. Optimal control theory and its applications: proceedings of the 14th biennial seminar of the Canadian Mathematical Congress, University of Western Ontario, Aug. 12–25, 1973, volume 106 of Lecture notes in economics and mathematical systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 3-540-07026-5. ISSN 0075-8442 (print), 2196-9957 (electronic). LCCN QA402.3 C33 1974.

Wang:1980:AIV

[199] Keith Y. Wang, editor. Acoustical Imaging: Visualization and Characterization, volume 9 of Acoustical Imaging. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 1-4684-3757-7, 1-4684-3755-0 (e-book). ISSN 0270-5117 (print), 2215-1869 (electronic). ???? pp. LCCN QC1-75. URL http://www.springerlink.

com/content/978-1-4684-3755-3.

Allgower:1981:NSN

[200] Eugene L. Allgower, Klaus Glashoff, and Heinz-Otto Peitgen, editors. Numerical Solution of Nonlinear Equations: Proceedings, Bremen 1980, volume 878 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNMAA2. ISBN 0-387-10871-8, 3-540-10871-8 (print), 3-540-38781-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L471

no.878; QA3.L471. URL http://link.springer.com/book/10.1007/
BFb0090674;   http://www.springerlink.com/content/978-3-540-
38781-7.

McAvoy:1983:IUS

[201] B. R. McAvoy, editor. IEEE 1983 Ultrasonics Symposium: October 31, November 1-2, 1983, Atlanta Marriott Hotel, Atlanta, Georgia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. CODEN ULSPDT. ISBN ???? ISSN 0090-5607. LCCN ???? Two volumes. IEEE catalog number 83CH1947-1.

Graves-Morris:1984:RAI

[202] Peter Russell Graves-Morris, Edward B. Saff, and Richard S. Varga, editors. Rational Approximation and Interpolation: Proceedings of the United Kingdom — United States Conference held at Tampa, Florida, December 12–16, 1983, volume 1105 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L471

no.1105; QA3.L471. URL http://link.springer.com/book/10.1007/
BFb0072395;   http://www.springerlink.com/content/978-3-540-
39113-5.

Kaveh:1984:AIP

[203] M. Kaveh, R. K. Mueller, and J. F. Greenleaf, editors. Acoustical imaging: Proceedings of the thirteenth International Symposium on Acoustical Imaging, held October 26–28, 1983, in Minneapolis, MN, volume 13 of Acoustical imaging. Plenum Press, New York, NY, USA; London, UK, 1984. CODEN ACIGD9. ISBN 0-306-41717-0. ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN ????

Miller:1984:CMA

[204] Anthony Miller, editor. Contributions of mathematical analysis to the numerical solution of partial differential equations, volume 7 of Proceedings of the Centre for Mathematical Analysis, Australian National University. Centre for Mathematical Analysis, Australian National University, Canberra, Australia, 1984. ISBN 0-86784-508-2. LCCN QA374 .C661 1984.

McAvoy:1985:IUS

[205] B. R. McAvoy, editor. IEEE 1985 Ultrasonics Symposium: October 16–18, 1985, Cathedral Hill Hotel, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. CODEN ULSPDT. ISBN ???? ISSN 0090-5607. LCCN TA367 U47 1985 v. 1-2. Two volumes. IEEE catalog number 85CH2209-5.

Nalcioglu:1986:IWP

[206] O. (Orhan) Nalcioglu, Z.-H. (Zang-Hee) Cho, Thomas F. (Thomas Francis) Budinger, et al., editors. International Workshop on Physics and Engineering of Computerized Multidimensional Imaging and Processing: 2–4 April 1986, Newport Beach, California, volume 671 of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 1986. CODEN PSISDG. ISBN 0-89252-706-4. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK8315 .I5451 1986; TK8315 .I57 1986; TA1632 .I62 1986; TS510 .S63; TA1632 .I58 1986.

Jones:1987:AIP

[207] Hugh W. Jones, editor. Acoustical Imaging: Proceedings of the International Symposium, July 14–16, 1986, Halifax, NS, Canada, volume 15 of Acoustical imaging. Plenum Press, New York, NY, USA; London, UK, 1987. CODEN ACIGD9. ISBN 0-306-42565-3. ISSN 0270-5117 (print), 2215-1869 (electronic). LCCN ????

Bowers:1989:CCP

[208] K. Bowers, J. Lund, K. L. (Kenneth L.) Bowers, and J. (John) Lund, editors. Computation and control: proceedings of the Bozeman conference, Bozeman, Montana, August 1–11, 1988, volume 1 of Progress in systems and control theory. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1989. ISBN 0-8176-3438-X. LCCN TA329 .C645 1989.

Martin:1990:VSL

[209] Clyde Martin and John White, editors. Visiting scholars’ lectures 1989, Texas Tech University, Lubbock, TX (USA), volume 16 of Mathematics Series. Department of Mathematics, Texas Tech University, Lubbock, TX, USA, 1990.

Wong:1990:ACA

[210] R. (Roderick) Wong, editor. Asymptotic and Computational Analysis: Conference in Honor of Frank W. J. Olver’s 65th Birthday, volume 124 of Lecture Notes in Pure and Applied Mathematics. Marcel Dekker, New York, NY, USA, 1990. ISBN 0-8247-8347-6. LCCN QA299.6

.A88 1990. URL http://www.loc.gov/catdir/enhancements/fy0647/
90002810-d.html.

Bowers:1991:CCI

[211] K. L. (Kenneth L.) Bowers and J. (John) Lund, editors. Computation and control II: proceedings of the second Bozeman conference, Bozeman, Montana, August 1–7, 1990, volume 11 of Progress in systems and control theory. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1991. ISBN 0-8176-3611-0. LCCN TA329 .C644 1991. US$65.00.

Espelid:1992:NIR

[212] Terje O. Espelid and Alan Genz, editors. Numerical integration: recent developments, software and applications, volume 357 of NATO ASI series.

Series C, Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1583-9. LCCN QA299.3 .N38 1991.

Stenger:1993:NMB

[213] Frank Stenger. Numerical methods based on Sinc and analytic functions, volume 20 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94008-1 (New York), 3-540-94008-1 (Berlin). xv + 565 pp. LCCN QA372 .S82 1993.

Zahar:1994:ACF

[214] R. V. M. (Ramsay Vincent Michael) Zahar, editor. Approximation and computation: a festschrift in honor of Walter Gautschi: proceedings of the Purdue conference, December 2–5, 1993, volume 119 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Ang:1995:IPA

[215] – D.ǎng – Diñh Áng et al., editors. Inverse problems and applications to geophysics, industry, medicine and technology: proceedings of the International Workshop on Inverse Problems, 17–19 January 1995, Ho Chi Minh City, volume 2 of Publications of the Ho Chi Minh City Mathematical Society. Vietnam Mathematical Society, Ho Chi Minh City, Vietnam, 1995. ISBN ???? LCCN ????

Ismail:1995:MAW

[216] Mourad Ismail et al., editors. Mathematical analysis, wavelets, and signal processing: an International Conference on Mathematical Analysis and Signal Processing, January 3–9, 1994, Cairo University, Cairo, Egypt, volume 190 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 1995. ISBN 0-8218-0384-0. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA299.6 .M38 1995.

Papamichael:1999:CMF

[217] N. (Nicolas) Papamichael, Stephan Ruscheweyh, and E. B. Saff, editors. Computational methods and function theory 1997: proceedings of the Third CMFT Conference, 13–17 October 1997, Nicosia, Cyprus, volume 11 of Series in approximations and decompositions. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. ISBN 981-02-3626-3. LCCN QA297 .I473 1997.

Kromann:2000:ISI

[218] Gary B. Kromann, J. Richard Culham, and Koneru Ramakrishna, editors. ITherm 2000: the Seventh Intersociety Conference on Thermal and Thermomechanical Phenomena in Electronic Systems, presented at Las Vegas, Nevada, USA, May 23–26, 2000. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7803-5912-7 (softcover), 0-7803-5913-5 (casebound), 0-7803-5914-3 (microfiche). LCCN TK7870.25 .I6 2000. Two volumes. IEEE catalog number 00CH37069.

Nashed:2002:IPI

[219] M. Zuhair Nashed and Otmar Scherzer, editors. Inverse problems, image analysis, and medical imaging: AMS Special Session on Interaction of Inverse Problems and Image Analysis, January 10–13, 2001, New Orleans, Louisiana, volume 313 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 2002. ISBN 0-8218-2979-3. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN TA1637 .A47 2001.

Shen:2013:MSA

[220] Xiaoping Shen and Ahmed I. Zayed, editors. Multiscale Signal Analysis and Modeling. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4144-7, 1-4614-4145-5 (e-book). xvii + 378 pp. LCCN TA342 .M855 2013.

Zayed:2014:NPA

[221] Ahmed I. Zayed and Gerhard Schmeisser, editors. New Perspectives on Approximation and Sampling Theory: Festschrift in Honor of Paul Butzer’s 85th Birthday. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-08800-9, 3-319-08801-7 (e-book). ISSN 2296-5009 (print), 2296-5017 (electronic). xxii + 472 pp. LCCN ????

Sabin:2015:CMP

[222] John R. Sabin and Remigio Cabrera-Trujillo, editors. Concepts of mathematical physics in chemistry: a tribute to Frank E. Harris, volume 71. Academic Press Inc., New York, USA, 2015. ISBN 0-12-802824-6, 0-12-802868-8 (e-book). xv + 382 pp. LCCN QC19.2.

Sabin:2016:CMP

[223] John R. Sabin and Remigio Cabrera-Trujillo, editors. Concepts of mathematical physics in chemistry: a tribute to Frank E. Harris. Part B, volume 72 of Advances in quantum chemistry. Academic Press Inc., New York, USA, 2016. ISBN 0-12-803984-1, 0-12-803985-X (e-book). ISSN 0065-3276 (print), 2162-8815 (electronic). xii + 235 pp. LCCN QC19.2.

Govil:2017:PAT

[224] Narendra Kumar Govil, Ram Mohapatra, Mohammed A. Qazi, and Gerhard Schmeisser, editors. Progress in Approximation Theory and Applicable Complex Analysis: In Memory of Q. I. Rahman, volume 117 of Springer Optimization and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-49240-3 (print), 3-319-49242-X (e-book). LCCN QA402.5-402.6. URL

https://link.springer.com/chapter/10.1007/978-3-319-49242-1.

Baumann:2021:NSM

[225] Gerd Baumann, editor. New Sinc methods of numerical analysis: Festschrift in honor of Frank Stenger’s 80th birthday. Trends in mathematics. Birkhäuser, Cham, Switzerland, 2021. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. xvi + 404 pp. LCCN QA372.